\begin{thebibliography}{10}

\bibitem{AB16stoc}
{\sc Abboud, A., and Bodwin, G.}
\newblock The 4/3 additive spanner exponent is tight.
\newblock In {\em Proceedings of the 48th Annual ACM SIGACT Symposium on Theory
  of Computing (STOC)\/} (2016), ACM Special Interest Group on Algorithms and
  Computation Theory, pp.~351--361.

\bibitem{ABP17}
{\sc Abboud, A., Bodwin, G., and Pettie, S.}
\newblock A hierarchy of lower bounds for sublinear additive spanners.
\newblock In {\em Proceedings of the 28th Annual ACM-SIAM Symposium on Discrete
  Algorithms (SODA)\/} (2017), Society for Industrial and Applied Mathematics,
  pp.~568--576.

\bibitem{ACIM99}
{\sc Aingworth, D., Chekuri, C., Indyk, P., and Motwani, R.}
\newblock Fast estimation of diameter and shortest paths (without matrix
  multiplication).
\newblock {\em SIAM Journal on Computing 28}, 4 (1999), 1167--1181.

\bibitem{ADDJS93}
{\sc Alth{\"o}fer, I., Das, G., Dobkin, D., Joseph, D., and Soares, J.}
\newblock On sparse spanners of weighted graphs.
\newblock {\em Discrete \& Computational Geometry 9}, 1 (1993), 81--100.

\bibitem{AFIR10}
{\sc Ausiello, G., Franciosa, P.~G., Italiano, G.~F., and Ribichini, A.}
\newblock Computing graph spanners in small memory: fault-tolerance and
  streaming.
\newblock {\em Discrete Mathematics, Algorithms and Applications 2}, 04 (2010),
  591--605.

\bibitem{BKMP10}
{\sc Baswana, S., Kavitha, T., Mehlhorn, K., and Pettie, S.}
\newblock Additive spanners and ($\alpha$, $\beta$)-spanners.
\newblock {\em ACM Transactions on Algorithms (TALG) 7}, 1 (2010), 5.

\bibitem{BK08}
{\sc Bernstein, A., and Karger, D.}
\newblock Improved distance sensitivity oracles via random sampling.
\newblock In {\em Proceedings of the nineteenth annual ACM-SIAM symposium on
  Discrete algorithms\/} (2008), Society for Industrial and Applied
  Mathematics, pp.~34--43.

\bibitem{BGLP16esa}
{\sc Bilo, D., Guala, L., Leucci, S., and Proietti, G.}
\newblock {Compact and Fast Sensitivity Oracles for Single-Source Distances}.
\newblock In {\em 24th Annual European Symposium on Algorithms (ESA 2016)\/}
  (Dagstuhl, Germany, 2016), P.~Sankowski and C.~Zaroliagis, Eds., vol.~57 of
  {\em Leibniz International Proceedings in Informatics (LIPIcs)}, Schloss
  Dagstuhl--Leibniz-Zentrum fuer Informatik, pp.~13:1--13:14.

\bibitem{BDPV18}
{\sc Bodwin, G., Dinitz, M., Parter, M., and Williams, V.~V.}
\newblock Optimal vertex fault tolerant spanners (for fixed stretch).
\newblock In {\em Proceedings of the 29th Annual ACM-SIAM Symposium on Discrete
  Algorithms (SODA)\/} (2018), Society for Industrial and Applied Mathematics,
  pp.~1884--1900.

\bibitem{BGPV17}
{\sc Bodwin, G., Grandoni, F., Parter, M., and Williams, V.~V.}
\newblock Preserving distances in very faulty graphs.
\newblock In {\em Proceedings of the 44th International Colloquium on Automata,
  Languages, and Programming (ICALP)\/} (2017), European Association for
  Theoretical Computer Science, pp.~73:1--73:14.

\bibitem{Chechik13soda}
{\sc Chechik, S.}
\newblock New additive spanners.
\newblock In {\em Proceedings of the 24th Annual ACM-SIAM Symposium on Discrete
  Algorithms (SODA)\/} (2013), SIAM, pp.~498--512.

\bibitem{CCFK17}
{\sc Chechik, S., Cohen, S., Fiat, A., and Kaplan, H.}
\newblock (1+ $\varepsilon$)-approximate f-sensitive distance oracles.
\newblock In {\em Proceedings of the Twenty-Eighth Annual ACM-SIAM Symposium on
  Discrete Algorithms\/} (2017), Society for Industrial and Applied
  Mathematics, pp.~1479--1496.

\bibitem{CLPR10}
{\sc Chechik, S., Langberg, M., Peleg, D., and Roditty, L.}
\newblock Fault tolerant spanners for general graphs.
\newblock {\em SIAM Journal on Computing 39}, 7 (2010), 3403--3423.

\bibitem{CZ04}
{\sc Czumaj, A., and Zhao, H.}
\newblock Fault-tolerant geometric spanners.
\newblock {\em Discrete \& Computational Geometry 32}, 2 (2004), 207--230.

\bibitem{DTCR08}
{\sc Demetrescu, C., Thorup, M., Chowdhury, R.~A., and Ramachandran, V.}
\newblock Oracles for distances avoiding a failed node or link.
\newblock {\em SIAM Journal on Computing 37}, 5 (2008), 1299--1318.

\bibitem{DK11podc}
{\sc Dinitz, M., and Krauthgamer, R.}
\newblock Fault-tolerant spanners: better and simpler.
\newblock In {\em Proceedings of the 30th annual ACM SIGACT-SIGOPS symposium on
  Principles of distributed computing\/} (2011), ACM, pp.~169--178.

\bibitem{DZ16}
{\sc Dinitz, M., and Zhang, Z.}
\newblock Approximating low-stretch spanners.
\newblock In {\em Proceedings of the twenty-seventh annual ACM-SIAM symposium
  on Discrete algorithms\/} (2016), Society for Industrial and Applied
  Mathematics, pp.~821--840.

\bibitem{DHZ00}
{\sc Dor, D., Halperin, S., and Zwick, U.}
\newblock All-pairs almost shortest paths.
\newblock {\em Siam Journal on Computing (SICOMP) 29}, 5 (2000), 1740--1759.

\bibitem{EP04}
{\sc Elkin, M., and Peleg, D.}
\newblock (1+$\varepsilon$,$\beta$)-spanner constructions for general graphs.
\newblock {\em SIAM Journal on Computing 33}, 3 (2004), 608--631.

\bibitem{girth}
{\sc Erd{\H{o}}s, P.}
\newblock Extremal problems in graph theory.
\newblock In {\em Proceedings of the Symposium on Theory of Graphs and its
  Applications\/} (1963), p.~2936.

\bibitem{FS16}
{\sc Filtser, A., and Solomon, S.}
\newblock The greedy spanner is existentially optimal.
\newblock In {\em Proceedings of the 2016 ACM Symposium on Principles of
  Distributed Computing\/} (2016), ACM, pp.~9--17.

\bibitem{GV12}
{\sc Grandoni, F., and Williams, V.~V.}
\newblock Improved distance sensitivity oracles via fast single-source
  replacement paths.
\newblock In {\em Foundations of Computer Science (FOCS), 2012 IEEE 53rd Annual
  Symposium on\/} (2012), IEEE, pp.~748--757.

\bibitem{LNS98}
{\sc Levcopoulos, C., Narasimhan, G., and Smid, M.}
\newblock Efficient algorithms for constructing fault-tolerant geometric
  spanners.
\newblock In {\em Proceedings of the thirtieth annual ACM symposium on Theory
  of computing\/} (1998), ACM, pp.~186--195.

\bibitem{Parter15}
{\sc Parter, M.}
\newblock Dual failure resilient bfs structure.
\newblock In {\em Proceedings of the 2015 ACM Symposium on Principles of
  Distributed Computing\/} (2015), ACM, pp.~481--490.

\bibitem{PP13}
{\sc Parter, M., and Peleg, D.}
\newblock Sparse fault-tolerant bfs trees.
\newblock In {\em European Symposium on Algorithms\/} (2013), Springer,
  pp.~779--790.

\bibitem{PP18}
{\sc Parter, M., and Peleg, D.}
\newblock Fault-tolerant approximate bfs structures.
\newblock {\em ACM Transactions on Algorithms (TALG) 14}, 1 (2018), 10.

\bibitem{PU89sicomp}
{\sc Peleg, D., and Ullman, J.}
\newblock An optimal synchronizer for the hypercube.
\newblock {\em SIAM Journal on Computing (SICOMP) 18}, 4 (1989), 740–--747.

\bibitem{PU89jacm}
{\sc Peleg, D., and Upfal, E.}
\newblock A trade-off between space and efficiency for routing tables.
\newblock {\em Journal of the ACM (JACM) 36}, 3 (1989), 510--530.

\bibitem{TZ06}
{\sc Thorup, M., and Zwick, U.}
\newblock Spanners and emulators with sublinear distance errors.
\newblock In {\em Proceedings of the 17th Annual ACM-SIAM Symposium on Discrete
  Algorithms (SODA)\/} (2006), Society for Industrial and Applied Mathematics,
  pp.~802--809.

\bibitem{WY13}
{\sc Weimann, O., and Yuster, R.}
\newblock Replacement paths and distance sensitivity oracles via fast matrix
  multiplication.
\newblock {\em ACM Transactions on Algorithms (TALG) 9}, 2 (2013), 14.

\end{thebibliography}

\documentclass{article}

\usepackage[utf8]{inputenc}
\usepackage[left=3cm, right=3cm, top=2cm, bottom=2cm]{geometry}

\usepackage{algorithm}
\usepackage[noend]{algpseudocode}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{authblk}

\usepackage{color}

\newtheorem{definition}{Definition}
\newtheorem{theorem}{Theorem}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}

\newcommand{\dist}{\text{dist}}

\newcommand{\Althofer}{Alth\"{o}fer}

\newcommand{\Erdos}{Erd\"{o}s}

\title{A Trivial Yet Optimal Solution to Vertex Fault Tolerant Spanners}
\author{Greg Bodwin\thanks{gbodwin6@cc.gatech.edu}\ }
\author{Shyamal Patel\thanks{shyamalpatelb@gmail.com}}
\affil{Georgia Tech}
\date{}

\begin{document}

\maketitle

\begin{abstract}
We give a short and easy upper bound on the worst-case size of fault tolerant spanners, which improves on all prior work and is fully optimal at least in the setting of vertex faults.
\end{abstract}

\section{Introduction}

This paper concerns \emph{spanners}, a fundamental primitive used in graph sketching.

\begin{definition} [Spanners]
A \emph{$k$-spanner} of a graph $G = (V, E, w)$ is a subgraph $H = (V, E' \subseteq E, w)$ for which
$\dist_H(s, t) \le k \cdot \dist_G(s, t)$ for all $s, t \in V$.
\end{definition}
\noindent Spanners have been intensively studied since the mid-80s \cite{PU89jacm, PU89sicomp, ADDJS93, ACIM99, DHZ00, EP04, TZ06, BKMP10, Chechik13soda, AB16stoc, FS16, ABP17}.
A staple in the literature is the \emph{greedy construction algorithm} of \Althofer{} et al \cite{ADDJS93}, which works as follows: initialize $H$ to be an empty graph, and then for each edge $(u, v)$ in the input graph $G$ in order of increasing weight, add $(u, v)$ to $H$ if currently $\dist_H(u, v) > k w_{(u, v)}$.
Besides its simplicity and obvious correctness, the greedy algorithm is popular because there is a simple proof that it is \emph{existentially optimal} \cite{ADDJS93, FS16}, meaning that the number of edges in $H$ never exceeds the worst-case number of edges needed for a $k$-spanner over all possible input graphs on as many nodes as $G$.

In practice, spanners are often applied to systems whose parts are prone to sporadic failures.
A spanner for such a system must be robust to these failures, giving rise to the notion of \emph{fault tolerance}:
\begin{definition} [Fault Tolerant Spanners]
Given a graph $G$, a subgraph $H$ is an \emph{$f$ Vertex Fault Tolerant (VFT), resp.\ Edge Fault Tolerant (EFT),} $k$-spanner of $G$ if for any set of $F$ vertices (resp.\ edges) in $G$ of size $|F| \le f$, $H \setminus F$ is a $k$-spanner of $G \setminus F$.
\end{definition}
\noindent To construct FT spanners, it is natural to consider the obvious adaptation of the greedy algorithm:

\begin{algorithm}
	\caption{The VFT (EFT) Greedy Algorithm}\label{greedy}
	\begin{algorithmic}
	   \Function{ft-greedy}{$G = (V, E, w), k, f$}
	   \State{$H \gets (V, \emptyset, w)$}
	   \For{ $(u,v) \in E$ in order of increasing weight}
	        \If{there exists a set $F$ of $|F| \le f$ vertices (edges) such that $\dist_{H \setminus F}(u, v) > k \cdot w(u,v)$}
	            \State{add $(u, v)$ to $H$}
	        \EndIf
	   \EndFor
       \State \Return{$H$}
       \EndFunction

	\end{algorithmic}
\end{algorithm}

Correctness is again obvious, but unfortunately the analyses used in the non-faulty setting all seem to break for the FT greedy algorithm.
Most prior work on FT spanners has thus abandoned the greedy approach in favor of more involved constructions \cite{LNS98, CZ04, CLPR10, DK11podc, AFIR10} (see also \cite{Parter15, PP13, PP18, BGPV17, CCFK17, DTCR08, BK08, GV12, WY13, DZ16, BGLP16esa}); an analysis of the FT greedy algorithm was only obtained recently via fairly complex arguments \cite{BDPV18}.
In contrast, our main result is a simple analysis of the FT greedy algorithm which improves on all of these previous bounds.

\begin{theorem} [Main Result] \label{thm:main}
Let $b(n,k)$ be the maximum possible number of edges in a graph on $n$ nodes and girth $>k$.
Then any graph $H$ on $n$ nodes returned by the VFT or EFT Greedy Algorithm with parameters $f, k$ satisfies
$$\left|E(H)\right| = O \left (f^2 \cdot b \left (\frac{n}{f}, k+1 \right) \right).$$
\end{theorem}
This upper bound is best possible in the VFT setting, for any construction algorithm, due to a simple lower bound construction in \cite{BDPV18} (meaning that any asymptotic tradeoff between $n, f, k, |E(H)|$ not promised by this theorem does not exist in general).
In the EFT setting, the bound of Theorem \ref{thm:main} was already known to be the best possible tradeoff when $k < 5$ \cite{BDPV18}, but for larger $k$ it is still conceivable to improve the upper bound as far as
$$\left|E(H)\right| \overset{?}{=} O\left(f \cdot b \left (\frac{n}{\sqrt{f}}, k+1\right) + nf\right).$$
It remains a major open question to asymptotically determine $b(n, k)$; the only known upper bound is the folklore \emph{Moore bounds} which state $b(n, k) = O\left(n^{1 + 1/\lfloor k/2 \rfloor}\right)$.
Plugging this into Theorem \ref{thm:main} yields:
\begin{corollary}
For any graph $H$ on $n$ nodes returned by the VFT or EFT Greedy Algorithm with parameters $f, 2k-1$, we have
$$\left|E(H)\right| = O \left( n^{1 + 1/k} f^{1 - 1/k} \right).$$
\end{corollary}
\noindent This corollary improves over the previous best upper bound in \cite{BDPV18} by a factor of $\exp(k)$.
The famous \emph{\Erdos{} girth conjecture} \cite{girth} posits that the Moore Bounds are tight, which would then imply that this corollary is best possible, at least for VFT spanners.

An open question left by this work is to improve the runtime of the FT greedy algorithm: in a naive implementation it is exponential in $f$.
It would be interesting to improve this dependence, or perhaps to find a different fast algorithm achieving the existential size bounds proved in this paper.
We note that \cite{DK11podc} gives a construction with polynomial runtime dependence on $f$, at the price of somewhat suboptimal spanner size.

\section{Proof of Main Result}

We will state the proof in the VFT setting here; the proof in the EFT setting is essentially identical.

\begin{definition} [Blocking Set] Given a graph $G = (V, E)$, a \emph{$k$-blocking set} for $G$ is a set $B \subseteq V \times E$ such that (1) every $(v,e) \in B$ has $v \not \in e$, and (2) for every cycle $C$ in $G$ on $\le k$ edges, there exists $(v,e) \in B$ such that $v,e \in C$.
\end{definition}

\begin{lemma} \label{lem:blocks}
Any graph $H$ returned by the VFT greedy algorithm with parameters $k, f$ has a $(k+1)$-blocking set of size at most $f|E(H)|$.
\end{lemma}
\begin{proof}
For each edge $e = (u,v) \in H$, let $F_e$ be the set of nodes such that $\dist_{H \setminus F_e}(u, v) > kw_e$ when $e$ is added to $H$.
Let
$$B := \left\{ \left(x, e \right) \ \mid \ e \in E(H), x \in F_e \right\};$$
since $|F_e| \le f$ for all $e$, we have $|B| \le f|E(H)|$.
We now show that $B$ is a $(k+1)$-blocking set.
Let $C$ be any cycle on $\le k+1$ edges in the final graph $H$ and let $e = (u,v)$ be the last edge in $C$ considered by the greedy algorithm.
By construction there is a $u \leadsto v$ path (through $C$) of total weight $\le k w_{(u,v)}$ when $e$ is added to $H$, and so some node $x \in C \setminus \{u, v\}$ must be included in $F_e$.
Thus $(x, e) \in B$.
\end{proof}

\begin{lemma} \label{lem:subgraph}
Let $H$ be any graph on $n$ nodes and $m$ edges, let $f = o(n)$ be a parameter, and suppose $H$ has a $(k+1)$-blocking set $B$ of size $|B| \le f |E(H)|$.
Then $H$ has a subgraph on $O(n/f)$ nodes, $\Omega(m/f^2)$ edges, and girth $> k+1$.
\end{lemma}
\begin{proof}

Let $H'$ be an induced subgraph of $H$ on a uniformly random subset of exactly $\lceil n/(2f) \rceil$ vertices, let $B' := B \cap \left(E(H') \times V(H')\right)$, and let $H''$ be obtained from $H'$ by deleting every edge contained in any pair in $B'$.
The graph $H''$ has girth $> k+1$, since by definition of blocking sets we have now deleted at least one node or edge from every cycle in $H$ on $\le k+1$ edges.
Additionally:
\begin{itemize}
\item Each edge in $E(H)$ survives in $E(H')$ iff both of its endpoints survive in $V(H')$, which happens with probability 
$$\frac{\left\lceil n/(2f) \right\rceil}{n} \cdot \frac{\left\lceil n/(2f) \right\rceil - 1}{n - 1} = (1+o(1)) \frac{1}{4f^2}.$$

\item Each pair $(x, (u, v)) \in B$ survives in $B'$ iff all of $x, u, v$ survive in $V(H')$, which happens with probability
$$\frac{\left\lceil n/(2f) \right\rceil}{n} \cdot \frac{\left\lceil n/(2f) \right\rceil - 1}{n - 1} \cdot \frac{\left\lceil n/(2f) \right\rceil - 2}{n - 2} = (1+o(1)) \frac{1}{8f^3} $$
\end{itemize}
We may now compute
$$\mathbb{E}\left[|E(H'')|\right] \ge \mathbb{E}\left[|E(H')| - |B'|\right] = (1+o(1)) \left( \frac{|E(H)|}{4f^2} - \frac{|B|}{8f^3} \right) \ge (1+o(1)) \left(\frac{m}{4f^2} - \frac{m}{8f^2} \right) = \Omega\left(\frac{m}{f^2}\right).$$
There exists a possible setting of $H''$ which matches or exceeds this expectation, which thus has $\Omega(m/f^2)$ edges and satisfies the lemma.
\end{proof}

\begin{proof} [Proof of Theorem \ref{thm:main}]
If $f = \Omega(n)$, then Theorem \ref{thm:main} holds trivially since it states $|E(H)| = O(n^2)$.
Otherwise, let $H$ be an output graph of the FT Greedy algorithm on $n$ edges and $m$ nodes.
By Lemmas \ref{lem:blocks} and \ref{lem:subgraph}, $H$ has a subgraph of girth $>k+1$ on $O(n/f)$ nodes and $\Omega(m/f^2)$ edges, and hence
\begin{align*}
b(O(n/f), k+1) &= \Omega\left(\frac{m}{f^2}\right)\\
m &= O\left(f^2 \cdot b(n/f, k+1)\right). \qedhere
\end{align*}
\end{proof}

We conclude by remarking on a limitation of our approach.
Our definition of blocking sets is very VFT-centric; since a gap remains in the EFT setting, it is tempting to try to adapt this definition to the EFT setting in search of better upper bounds.
In particular, let us say that an \emph{edge $k$-blocking set} is a set of distinct edge pairs such that every $(\le k)$-cycle $C$ has $e_1, e_2 \in C$ for some $(e_1, e_2) \in B$.
It is easy to show that any graph $H$ returned by the EFT greedy algorithm with parameters $f, k$ admits an edge $(k+1)$-blocking set of size $\le f|E(H)|$ (the analog of Lemma \ref{lem:blocks}).
We would then need an \emph{improved} analog of Lemma \ref{lem:subgraph} in order to get improved upper bounds on EFT spanners for $k \ge 5$.
However, no such improvement is possible: for any $k$ we can show a graph $H$ on $\Omega(f^2 \cdot b(n/f, k+1))$ edges that has an edge $(k+1)$-blocking set of size $\le f|E(H)|$, and so our analog of Lemma \ref{lem:blocks} alone is not powerful enough to get improved upper bounds in the EFT setting.
Specifically, this $H$ is the same as the VFT lower bound graph of \cite{BDPV18}: it is the Cartesian product of an arbitrary graph of girth $> k+1$ with a biclique on $\lfloor f/2 \rfloor$ nodes; the blocking set is then all pairs of edges that share an endpoint in the product graph and which correspond to the same edge in the initial high-girth graph.
Hence any improvement to our EFT upper bounds (if possible) will need to exploit stronger properties of $H$ than just the existence of a small edge blocking set.

\bibliographystyle{acm}
\bibliography{References}

\end{document}